OpenStudy (anonymous):

2 + a/-4 = 3/5 Please solve and show steps. Thanks!!

OpenStudy (anonymous):

i would start by multiplying everything by -4, and then by 5. this will get rid of the denominators: -8+a = -12/5 -40+5a = -12 can you solve from there?

OpenStudy (anonymous):

Ya, thanks!

OpenStudy (anonymous):

np